OpenStudy (anonymous):

If 50 g of liquid water [C = 4,180 J/(kg*K)] changes temperature in a calorimeter from 298 K to 348 K, then how much heat is transferred? A. 10.5 J B. 10.5 kJ C. 72.7 J D. 100 kJ **Not sure about this one! thank you:)

OpenStudy (michele_laino):

here we have to apply this formula: \[\Large Q = cm\Delta T = 4180 \times 0.05 \times \left( {348 - 298} \right) = ...Joules\]

OpenStudy (anonymous):

okay!! so we get 10450 Joules? so about 10.5 kJ ?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay!! :D thanks!!! :D

OpenStudy (michele_laino):

thanks!!!!! :D